Find the integral of x2 dx dy. Use inner limits = 0 to 2y and outer limits = 0 to 2. (5 points)

1
Expert's answer
2022-05-24T09:08:01-0400
"\displaystyle\int_{0}^{2}\displaystyle\int_{0}^{2y}x^2dxdy=\displaystyle\int_{0}^2[\dfrac{x^3}{3}]\begin{matrix}
 2y \\
 0
\end{matrix}dy"

"=\displaystyle\int_{0}^2\dfrac{8y^3}{3}dy=[\dfrac{2y^4}{3}]\begin{matrix}
2 \\
0
\end{matrix}=\dfrac{32}{3}"